OpenStudy (anonymous):

1.log 20 2.In 14 3.log3 10 (And what does In stand for?)

OpenStudy (misty1212):

what is the question ?

OpenStudy (anonymous):

ln stands for natural log e

OpenStudy (solomonzelman):

\(\large\color{slate}{ \ln }\) is a natural log. it is log base e, For example, when you say: \(\large\color{slate}{ \ln(4) }\) it is same thing as, \(\large\color{slate}{ \log_e(4) }\). \(\large\color{blue}{ \ln}\) is same as \(\large\color{blue}{ \log_e}\)